OpenStudy (anonymous):

A zoologist is recording the loss of deer in a forest. She notes the number of deer, d, in the forest on January 1. One year later, there were 73 deer in the forest, which is 19 fewer deer than were in the forest a year earlier. Which equation represents this situation? A. d + 19 = 73 B. d – 73 = –19 C. 73 – d = 19 D. d – 19 = 73

OpenStudy (anonymous):

A or C @Vocaloid

OpenStudy (anonymous):

@mrqueest

OpenStudy (anonymous):

I would say C

OpenStudy (anonymous):

ill choose it and see what i get.

OpenStudy (anonymous):

On Wednesday, William wrote a check in the amount of $45.60. This brought his balance down to –$15.75. Suppose n represents William’s balance before he wrote the check. Which equation can be used to represent this situation? A. n + 15.75 = –45.60 B. n – 15.75 = 45.60 C. n + 45.60 = –15.75 D. n – 45.60 = –15.75

OpenStudy (anonymous):

I think D

OpenStudy (anonymous):

Yeah I would say that too

OpenStudy (anonymous):

the first one the answer was D

OpenStudy (anonymous):

I have to retake it

OpenStudy (anonymous):

sorry about that lol not the best with word problems, at least you get to retake it :)

OpenStudy (anonymous):

I have 2 retries !:D its okay

OpenStudy (anonymous):

so lucky, I wish i had that for my ap stats class, so I could actually pass a quiz haha

OpenStudy (anonymous):

Haha!!!

OpenStudy (anonymous):

This morning, Elena wrote a check in the amount of $42.50. This brought her balance down to –$14.25. Suppose n represents Elena's balance before she wrote the check. Which equation can be used to represent this situation? A. n – 14.25 = 42.50 B. n + 42.50 = –14.25 C. n – 42.50 = –14.25 D. n – 14.25 = –42.50